OpenStudy (anonymous):

work needing to be checked thus far: f(x)=x^3 and g(x) =5+4x Find (fog)(X) my work: (fog)(x) = f(g(x)) =f(5+4x)^3 im not sure if that's right to start out with ...

OpenStudy (anonymous):

Its right.

OpenStudy (bibby):

(fog)(x) = f(g(x)) is correct. that's how the composite function works you replace the x in f(x) with (5+4x)

OpenStudy (anonymous):

okay I wasn't sure if I placed the ^3 in the correct place

OpenStudy (anonymous):

ty is my answer 64x^3 + 125?

OpenStudy (bibby):

I dunno is it? :) Post your work as I'm gettin ga different answer

OpenStudy (anonymous):

yes! =)

OpenStudy (bibby):

\[64x^3+240x^2+300x+15\]

OpenStudy (anonymous):

(fog)(x)=f(5+4x)^3 =125 +64x^3

OpenStudy (anonymous):

the one u got is very not right lol

OpenStudy (bibby):

I checked wolfram alpha .com. I'm pretty sure I'm right

OpenStudy (anonymous):

that's not what any of my answers look like

OpenStudy (bibby):

I actually multiplied out (5+4x)(5+4x)(5+4x) http://www.wolframalpha.com/input/?i=%285%2B4x%29%5E3 I'm almost positive I'm right and you can call someone else to check. Either way (5+4x)^3 should be fine as an answer

OpenStudy (anonymous):

thats what one of my answers look like . oh sorry I didn't even think about going as far as you did ... your good!

OpenStudy (bibby):

Sorry if I come across as a jerk. I'm glad I could help in any sort of manner